# Thermal Area Law in Long-Range Interacting Systems

### v2

@@ Abstract
where $h_{Z}$ is the local Hamiltonian on a subset $Z\subset\Lambda$ with site number $|Z|$. We consider the long-range interactions satisfying the following condition:
@@
where $d_{X,Y}=\min_{i\in X,j\in Y}d_{i,j}$ and $|X|$ and $|Y|$ are numbers of sites in the regions $X$ and $Y$, respectively. $C$ is an $\mathcal{O}(1)$ constant and the threshold temperature is given by $\beta_{c}=1/(8ugk)$ with $u=\sum_{j\in\Lambda}1/(1+d_{i,j})^{\alpha}$.
@@
We first verify the clustering property (6) in the Hamiltonian (9[55, 56]. We analyze a 1D chain of $N=1000$ sites and a 2D square lattice of side length $N=40$, focusing on two-point correlation functions $\langle c_{i}^{\dagger}c_{i+r}\rangle$ in 1D and $\langle c_{{\bm{i}}}^{\dagger}c_{{\bm{i}}+{\bm{r}}}\rangle$ in 2D at $\beta=2$. For general bilinear systems, we assign random values $\in[0,1]$ to the hopping variable and average over 1000 samples. In Figs. 2(a) and (b), $|\langle c_{i}^{\dagger}c_{i+r}\rangle|\times r^{\alpha}$ (1D) and $|\langle c_{{\bm{i}}}^{\dagger}c_{{\bm{i}}+{\bm{r}}}\rangle|\times|{\bm{r}}|^{\alpha}$ (2D) saturate at large distances for any $\alpha$, confirming clustering in both dimensions, even for $\alpha<D$. While we show the averaged data, we checked that individual data also satisfy the clustering property (not shown). All correlation functions in bilinear systems follow this property via the Wick theorem.
@@
It can be proven that the summation of the distance $\sum_{i\in A}\sum_{j\in B}d_{i,j}^{-2\alpha}$ is upper bounded by the boundary area when $12\alpha>D+$ 1. Therefore, inequality (19) leads to the desired inequality (7).
@@ acknowledgments
We consider a quantum system, where we use $\Lambda$ to represent the set of sites. Given an arbitrary subset $X$ in $\Lambda$, $X\subseteq\Lambda$, we denote the cardinality of the set $X$, i.e., the number of sites contained in $X$, by $|X|$. For any two subsets $X,Y$ in $\Lambda$, $X,Y\subseteq\Lambda$, we define the distance between the two sets, $d_{X,Y}$, as the shortest path length of the graph that connects $X$ and $Y$. We remark that when $X$ intersects with $Y$, $X\cap Y\neq\emptyset$, the distance between the sets $X$ and $Y$ is zero: $d_{X,Y}=0$. We denote the complementary set of $X$ as $X^{\rm c}$, namely $X^{\rm c}:=\Lambda\setminus X$, and the surface subset of $X$ as $\partial X$, i.e. $\partial X:=\{i\in X|d_{i,X^{\rm c}}=1\}$, respectively. Also, for an arbitrary operator $O$, we denote the support of $O$ by ${\rm Supp}(O)$. We often describe the support of the operator explicitly by adding an Index of lower right subscript such as $O_{X}$ with ${\rm Supp}(O_{X})=X$.
@@
where $Z$ is a subset indicating the interacting sites, and $h_{Z}$ is the local Hamiltonian. Note that $|Z|$ is the number of the interacting sites, and hence $h_{Z}$ means $Z$-body interaction. We assume that the interaction form satisfies the following property
@@
It was proven that $\sum_{i\in A}\sum_{j\in B}d_{i,j}^{-2\alpha}$ is upper bounded by the boundary area $|\partial A|$ when $12\alpha>D+$ 1 [39]. With Eq. (S.19), we finally obtain
@@
$\left\|\widehat{H_{L}}\right\|\leq g|L|,\quad\widehat{H_{L}}=\sum_{Z:Z\cap L\neq\emptyset}h_{Z}=H-H_{L^{\rm c}}$ (S.21)
for an arbitrary subset $L\subseteq\Lambda$. Then, the clustering theorem for the bipartite correlations is rigorously proven as
@@
where $C$ is an $\mathcal{O}(1)$ constant and the threshold temperature is given by $\beta_{c}=W(1/ue)/(2gk)$ with the Lambert $W$ function (i.e., the inverse function of $xe^{x}$) and $u=\sum_{j\in\Lambda}1/(1+d_{i,j})^{\alpha}$. We can lower-bound $\beta_{c}$ as $\beta_{c}\geq 1/(8ugk)$ by using
$W(1/ue)\leq 1/(1+ue)\leq 1/(4u),$
where we use $u$ is always larger than $1$ from the definition.
@@
$\|{\rho_{{\rm cl}}}\|_{1}$ $\leq{1\over{\cal Z}^{2}}\sum_{s=0}^{\infty}{\beta^{s}\over s!}\sum_{w_{\rm cl}\in\mathcal{G}_{\rm cl}}\|\tilde{h}(Z_{1},\cdots,Z_{s})\|\,{\rm tr}\Bigl{[}e^{-\beta H^{(+)}_{V_{w^{\rm c}_{\rm cl}}}}\Bigr{]}\leq\sum_{s=0}^{\infty}{\beta^{s}\over s!}\sum_{w_{\rm cl}\in\mathcal{G}_{\rm cl}}e^{2\beta g|V_{w_{\rm cl}}|}\|\tilde{h}(Z_{1},\cdots,Z_{s})\|\,.$ (S.37)
@@
$\|\rho_{{\rm cl}}\|_{1}$ $\leq\sum_{s=0}^{\infty}{(\beta e^{2\beta gk})^{s}\over s!}\sum_{s^{\prime}=1}^{s}{s!\over s^{\prime}!(s-s^{\prime})!}\sum_{\{Z_{1},\cdots,Z_{s^{\prime}}\}=w_{\rm cl}^{\rm s}}\|\tilde{h}(Z_{1},\cdots,Z_{s^{\prime}})\|^{\prime}\sum_{\{Z_{s^{\prime}+1},\cdots,Z_{s}\}=w_{\rm cl}^{\rm r}}\|\tilde{h}(Z_{s^{\prime}+1},\cdots,Z_{s})\|^{\prime}$
$=\sum_{s=0}^{\infty}(\beta e^{2\beta gk})^{s}\sum_{s^{\prime}=1}^{s}\sum_{w_{\rm cl}^{\rm s}}{\|\tilde{h}(Z_{1},\cdots,Z_{s^{\prime}})\|^{\prime}\over s^{\prime}!}\sum_{w_{\rm cl}^{\rm r}}{\|\tilde{h}(Z_{s^{\prime}+1},\cdots,Z_{s})\|^{\prime}\over(s-s^{\prime})!}\,,$ (S.39)
@@
We next consider the terms of $w_{\rm cl}^{\rm r}$ in (S.39). We define $n:=s-s^{\prime}$ and $L:=X\cup Y\cup\omega_{\rm cl}^{\rm s}$ for a simplicity. In addition, let us denote $Z_{s+m}$ by $Z^{\prime}_{m}\,(m=1,\cdots,n)$ for notational simplicity.
${1\over n!}\sum_{w_{\rm cl}^{\rm r}}\|\tilde{h}(Z^{\prime}_{1}\cdots,Z_{n}^{\prime})\|^{\prime}$ $={2^{n}\over n!}\bigl{(}\prod_{m=1}^{n}\sum_{Z^{\prime}_{m}}\|h_{Z^{\prime}_{m}}\|\bigr{)}\Bigr{|}_{\{L,Z^{\prime}_{1},\cdots,Z^{\prime}_{n}\}:{\rm nonsep}}\,$
$\leq{2^{n}\over n!}\sup_{Z^{\prime\prime}_{n}:|Z^{\prime\prime}_{n}|\leq k}\bigl{(}\prod_{m=1}^{n-1}\sum_{Z^{\prime}_{m}}\|h_{Z^{\prime}_{m}}\|\bigr{)}\Bigr{|}_{\{L,Z^{\prime}_{1},\cdots,Z^{\prime}_{n-1},Z^{\prime\prime}_{n}\}:{\rm nonsep}}\sum_{Z^{\prime}_{n}:Z^{\prime}_{n}\cap\{L,Z^{\prime}_{1},\cdots,Z^{\prime}_{n-1},Z^{\prime\prime}_{n}\}\neq\emptyset}\|h_{Z^{\prime}_{n}}\|\,$
$\leq{2^{n}\over n!}\sup_{Z^{\prime\prime}_{n}:|Z^{\prime\prime}_{n}|\leq k}\bigl{(}\prod_{m=1}^{n-1}\sum_{Z^{\prime}_{m}}\|h_{Z^{\prime}_{m}}\|\bigr{)}\Bigr{|}_{\{L,Z^{\prime}_{1},\cdots,Z^{\prime}_{n-1},Z^{\prime\prime}_{n}\}:{\rm nonsep}}gu(|L|+kn)\,,$ (S.45)
where |{$|_{\{A\}:{\rm nonsep}}$ implies taking a “nonseparable” set $\{A\}$, defined as a set where every subset can be reached from any starting subset by successively traversing overlapping subsets. We also use the following inequality to get the last line in (S.45):
$\sum_{Z^{\prime}_{n}:Z^{\prime}_{n}\cap\{L,Z^{\prime}_{1},\cdots,Z^{\prime}_{n-1},Z^{\prime\prime}_{n}\}\neq\emptyset}\|h_{Z^{\prime}_{n}}\|$ $\leq\sum_{i\in\{L,Z^{\prime}_{1},\cdots,Z^{\prime}_{n-1},Z^{\prime\prime}_{n}\}}\sum_{j\in\Lambda}\sum_{Z^{\prime}_{n}\ni(i,j)}\|h_{Z^{\prime}_{n}}\|\leq gu(|L|+kn)\,.$ (S.46)
We perform the same calculation as in (S.45) and (S.46) for the part in (S.45):
$\bigl{(}\prod_{m=1}^{n-1}\sum_{Z^{\prime}_{m}}\|h_{Z^{\prime}_{m}}\|\bigr{)}\Bigr{|}_{\{L,Z^{\prime}_{1},\cdots,Z^{\prime}_{n-1},Z^{\prime\prime}_{n}\}:{\rm nonsep}}$
$\leq\sup_{Z^{\prime\prime}_{n-1},Z^{\prime\prime}_{n}\atop:|Z^{\prime\prime}_{n-1}|\leq k,|Z^{\prime\prime}_{n}|\leq k}\bigl{(}\prod_{m=1}^{n-2}\sum_{Z^{\prime}_{m}}\|h_{Z^{\prime}_{m}}\|\bigr{)}\Bigr{|}_{\{L,Z^{\prime}_{1},\cdots,Z^{\prime}_{n-2},Z^{\prime\prime}_{n-1},Z^{\prime\prime}_{n}\}:{\rm nonsep}}\sum_{Z^{\prime}_{n-1}:Z^{\prime}_{n-1}\cap\{L,Z^{\prime}_{1},\cdots,Z^{\prime}_{n-2},Z^{\prime\prime}_{n-1},Z^{\prime\prime}_{n}\}\neq\emptyset}\|h_{Z^{\prime}_{n-1}}\|\,$
$\leq\sup_{Z^{\prime\prime}_{n-1},Z^{\prime\prime}_{n}\atop:|Z^{\prime\prime}_{n-1}|\leq k,|Z^{\prime\prime}_{n}|\leq k}\bigl{(}\prod_{m=1}^{n-2}\sum_{Z^{\prime}_{m}}\|h_{Z^{\prime}_{m}}\|\bigr{)}\Bigr{|}_{\{L,Z^{\prime}_{1},\cdots,Z^{\prime}_{n-2},Z^{\prime\prime}_{n-1},Z^{\prime\prime}_{n}\}:{\rm nonsep}}gu(|L|+kn)\,.$ (S.47)
By the iterative use of these computations, we obtain the following bound
${1\over n!}\sum_{w_{\rm cl}^{\rm r}}\|\tilde{h}(Z_{s^{\prime}+1}\cdots,Z_{s})\|^{\prime}$ $\leq{(2gu)^{n}\over n!}(|L|+nk)^{n}$
$\leq(2guek)^{n}e^{|L|/k}\leq(2guek)^{n}e^{(|X|+|Y|+s^{\prime}k)/k}\,,$ (S.48)
@@
where we have used the relations, $(n/e)^{n}<n!$, $(1+x/n)^{n}<e^{x}$ and $|L|<|X|+|Y|+s^{\prime}k$. Finally we sum over $s^{\prime}$ to get the following bound.
$\|\rho_{\rm cl}\|_{1}$ $\leq u^{-1}|X|\,|Y|{e^{(|X|+|Y|)/k}\over(1+d_{X,Y})^{\alpha}}\sum_{s=0}^{\infty}(2\beta gkeu\,e^{2\beta gk})^{s}\sum_{s^{\prime}=1}^{s}{(1/k)^{s^{\prime}}}$
$\leq(u(k-1))^{-1}|X|\,|Y|{e^{(|X|+|Y|)/k}\over(1+d_{X,Y})^{\alpha}}{1\over 1-2\beta gkeu\,e^{2\beta gk}}\,.$ (S.49)
@@
Fig. 6(a) illustrates the average mutual information computed over 20,000 random values of $t_{i,j}$ sampled from the interval $[0,1]$. The results reveal that the mutual information increases following the behavior of $N^{2-2\alpha}$ for $\alpha<1$, while saturating when $\alpha>1$.
@@
Figure 6: (a) The mutual information $\mathcal{I}_{\rho_{\beta}}(A:B)$ between subsystem $A$ (the first $N/2$ sites) and subsystem $B$ (the remaining $N/2$ sites) is computed. The system size $N$ varies from 100 to 1000 in increments of 100, while the interaction power $\alpha$ varies from 0.5 to 1.4 in increments of 0.1. The data represents the average mutual information over 20,000 samples, with the hopping parameters $t_{i,j}$ randomly selected from the interval $[0,1]$. For $\alpha<1$, a line proportional to $N^{2-2\alpha}$ is included in the figure for comparison. (b) The average mutual information at $\alpha=1.0$, calculated from 100,000 samples, is plotted as circles with the $x$-axis representing the natural logarithm of the system size, $\ln N$. The red line corresponds to the linear fit, given by $\mathcal{I}_{\rho_{\beta}}(A:B)=0.0575\ln N+0.2082$. (c) The circles (diamonds) represent the exponents $\Delta$, obtained by fitting the average mutual information from panel (a) to a power-law, $\mathcal{I}_{\rho_{\beta}}(A:B)\approx cN^{\Delta}$, over the ranges $N=100$ to $N=500$ (resp. $N=600$ to $N=1000$). The red line represents $\Delta=2-2\alpha$. For $\alpha<1$, the exponent $\Delta$ closely follows the line $2-2\alpha$, whereas for $\alpha>1$, $\Delta$ stays near zero. This behavior becomes more pronounced as the system size increases.
@@
where $a_{i,j}$ is a parameter of order $\mathcal{O}(1)$, $N$ is the number of sites, and ${\bm{S}}_{i}=(S_{i}^{x},S_{i}^{y},S_{i}^{z})$ is a spin-1/2 operator at site $i$. To support our argument that the thermal area holds for $\alpha>(D+1)/2$ and this criterion is optimal, we compute the mutual information $\mathcal{I}_{\rho_{\beta}}(A:B)$ between subsystems $A$ and $B$ at $\beta=2$, with $A$ representing the first $N/2$ sites and $B$ the remaining $N/2$ sites, by using the exact diagonalization method. We randomly assign $a_{i,j}$ from the interval $[0,1]$, and then compute the average over 10000 samples. Fig. 7 shows that the mutual information increases with the system size when $\alpha<1$, while it does not exhibit an increase for $\alpha>1$.
Figure 7: Exact diagonalization results: the averaged mutual information $\mathcal{I}_{\rho_{\beta}}(A:B)$ over 10000 samples of the long-range Heisenberg chain (S.51) with random numbers $a_{i,j}$ assigned from the interval $[0,1]$. Here $N$ is the number of sites, $\beta=2$, and $A$ and $B$ are the first $N/2$ sites and the remaining $N/2$ sites, respectively.
@@
We consider $n$ spinless fermionic particles with the basis states $|0\rangle$ and $|1\rangle=c^{\dagger}|0\rangle$. The determinant of the $n\times n$ matrix $\mathbf{X}=(x_{i,j})$ is defined as

### v3

@@ Abstract
[…]
@@
[…]
@@
We first verify the clustering property (6) in the Hamiltonian (9[55, 56]. We analyze a 1D chain of $N=1000$ sites and a 2D square lattice of side length $N=40$, focusing on two-point correlation functions $\langle c_{i}^{\dagger}c_{i+r}\rangle$ in 1D and $\langle c_{{\bm{i}}}^{\dagger}c_{{\bm{i}}+{\bm{r}}}\rangle$ in 2D at $\beta=2$. For general bilinear systems, we assign random values [0,1$\in$[0,1]∈ [ 0 , 1 ] to the hopping variable and average over 1000 samples. In Figs. 2(a) and (b), $|\langle c_{i}^{\dagger}c_{i+r}\rangle|\times r^{\alpha}$ (1D) and $|\langle c_{{\bm{i}}}^{\dagger}c_{{\bm{i}}+{\bm{r}}}\rangle|\times|{\bm{r}}|^{\alpha}$ (2D) saturate at large distances for any $\alpha$, confirming clustering in both dimensions, even for $\alpha<D$. While we show the averaged data, we checked that individual data also satisfy the clustering property (not shown). All correlation functions in bilinear systems follow this property via the Wick theorem.
@@
It can be proven that the summation of the distance $\sum_{i\in A}\sum_{j\in B}d_{i,j}^{-2\alpha}$ is upper bounded by the boundary area when $2\alpha>D+1$. Therefore, inequality (19) leads to the desired inequality (7).
@@ acknowledgments
[…]
@@
[…]
@@
It was proven that $\sum_{i\in A}\sum_{j\in B}d_{i,j}^{-2\alpha}$ is upper bounded by the boundary area $|\partial A|$ when $2\alpha>D+1$ [39]. With Eq. (S.19), we finally obtain
@@
$\widehat{H_{L}}$ $:=\sum_{Z:Z\cap L\neq\emptyset}h_{Z}=H-H_{L^{\rm c}}\,,$ (S.21)
$\left\|\widehat{H_{L}}\right\|$ $\leq\sum_{i\in L}\sum_{Z\in i}\|h_{Z}\|\leq gu|L|\,,$ (S.22)
for an arbitrary subset $L\subseteq\Lambda$. Here $u:=2^{\alpha}\sum_{j\in\Lambda}1/(1+d_{i,j})^{\alpha}$. Then, the clustering theorem for the bipartite correlations is rigorously proven as
@@
where $C$ is an $\mathcal{O}(1)$ constant and the threshold temperature is given by $\beta_{c}=W(1/2e^{2})/(2guk)$ with the Lambert $W$ function (i.e., the inverse function of $xe^{x}$) and $u=2^{\alpha}\sum_{j\in\Lambda}1/(1+d_{i,j})^{\alpha}$. We can lower-bound $\beta_{c}$ as $\beta_{c}\geq 1/(16guek)$ with the relation $W(1/2e^{2})>1/(8e)$.
@@
$\|{\rho_{{\rm cl}}}\|_{1}$ $\leq{1\over{\cal Z}^{2}}\sum_{s=0}^{\infty}{\beta^{s}\over s!}\sum_{w_{\rm cl}\in\mathcal{G}_{\rm cl}}\|\tilde{h}(Z_{1},\cdots,Z_{s})\|\,{\rm tr}\Bigl{[}e^{-\beta H^{(+)}_{V_{w^{\rm c}_{\rm cl}}}}\Bigr{]}\leq\sum_{s=0}^{\infty}{\beta^{s}\over s!}\sum_{w_{\rm cl}\in\mathcal{G}_{\rm cl}}e^{2\beta gu|V_{w_{\rm cl}}|}\|\tilde{h}(Z_{1},\cdots,Z_{s})\|\,.$ (S.38)
@@
$\|\rho_{{\rm cl}}\|_{1}$ $\leq\sum_{s=0}^{\infty}{(\beta e^{2\beta guk})^{s}\over s!}\sum_{s^{\prime}=1}^{s}{s!\over s^{\prime}!(s-s^{\prime})!}\sum_{\{Z_{1},\cdots,Z_{s^{\prime}}\}=w_{\rm cl}^{\rm s}}\|\tilde{h}(Z_{1},\cdots,Z_{s^{\prime}})\|^{\prime}\sum_{\{Z_{s^{\prime}+1},\cdots,Z_{s}\}=w_{\rm cl}^{\rm r}}\|\tilde{h}(Z_{s^{\prime}+1},\cdots,Z_{s})\|^{\prime}$
$=\sum_{s=0}^{\infty}(\beta e^{2\beta guk})^{s}\sum_{s^{\prime}=1}^{s}\sum_{w_{\rm cl}^{\rm s}}{\|\tilde{h}(Z_{1},\cdots,Z_{s^{\prime}})\|^{\prime}\over s^{\prime}!}\sum_{w_{\rm cl}^{\rm r}}{\|\tilde{h}(Z_{s^{\prime}+1},\cdots,Z_{s})\|^{\prime}\over(s-s^{\prime})!}\,,$ (S.40)
@@
We next consider the terms of $w_{\rm cl}^{\rm r}$ in (S.40). We define $n:=s-s^{\prime}$ and $L:=X\cup Y\cup\omega_{\rm cl}^{\rm s}$ for a simplicity. In addition, let us denote $Z_{s+m}$ by $\xi_{m}\,(m=1,\cdots,n)$ for notational simplicity. For a given set $\xi_{L}:=\{\xi_{m}\}_{m=1}^{n}$, we decompose it as follows.
$\xi_{L}$ $=\xi^{(1)}\oplus\xi^{(2)}\oplus\cdots\oplus\xi^{(\ell)}\,,~~~~~1\leq\ell\leq n\,$ (S.46)
where the set $\xi^{(j)}$ has the overlap $\xi^{(j-1)}$ with setting $\xi^{(0)}:=L$. Hence, $\xi^{(j)}$ is a achievable set from $L$ by the $j$-step connection. See the schematic in Fig.6. Let $n_{j}=|\xi^{(j)}|$ where $n_{j}\geq 1$. We take account of all possible combination of this structure:
$\sum_{w_{\rm cl}^{\rm r}}{\|\tilde{h}(\xi_{1},\cdots,\xi_{n})\|\over n!}$ $={2^{n}\over n!}\sum_{(\xi_{1},\cdots,\xi_{n})}\|h_{\xi_{1}}\|\cdots\|h_{\xi_{n}}\|$
$={2^{n}\over n!}\sum_{\ell=1}^{n}\sum_{n_{1}+\cdots+n_{\ell}=n,n_{i}\geq 1}{n!\over n_{1}!\cdots n_{\ell}!}$
×$\times\left(\sum_{\xi^{(1)}_{i}\in\xi^{(1)}|L}\|h_{\xi^{(1)}_{i}}\|\right)^{n_{1}}\left(\sum_{\xi^{(2)}_{i}\in\xi^{(2)}|\{L,\xi^{(1)}\}}\|h_{\xi^{(2)}_{i}}\|\right)^{n_{2}}\cdots\left(\sum_{\xi^{(\ell)}_{i}\in\xi^{(\ell)}|\{L,\xi^{(1)},\cdots,\xi^{(\ell-1)}\}}\|h_{\xi^{(\ell)}_{i}}\|\right)^{n_{\ell}}\,.$ (S.47)
Here, $\xi^{(j)}|\{L,\xi^{(1)},\cdots,\xi^{(j-1)}\}$ implies taking a set $\xi^{(j)}$ for fixed set $\{L,\xi^{(1)},\cdots,\xi^{(j-1)}\}$. We first upper-bound the most right term in (S.47):
$\sum_{\xi^{(\ell)}_{i}\in\xi^{(\ell)}|\{L,\xi^{(1)},\cdots,\xi^{(\ell-1)}\}}\|h_{\xi^{(\ell)}_{i}}\|$ $\leq\sum_{j\in\{\xi^{(\ell-1)}\}}\sum_{\xi^{(\ell)}_{i}\ni j}\|h_{\xi^{(\ell)}_{i}}\|\leq(kn_{\ell-1})gu\,,$ (S.48)
where we use $|\xi^{(\ell-1)}|=n_{\ell-1}$ and $\sum_{\xi\ni i}\|h_{\xi}\|\leq gu$. Iterative use of this computation down to the summation over $\xi_{i}^{(1)}$, we obtain the following bound
$\sum_{w_{\rm cl}^{\rm r}}{\|\tilde{h}(\xi_{1},\cdots,\xi_{n})\|\over n!}$ $\leq{(2kgu)^{n}}\sum_{\ell=1}^{n}\sum_{n_{1}+\cdots+n_{\ell}=n\atop n_{i}\geq 1}{1\over n_{1}!\cdots n_{\ell}!}(|L|/k)^{n_{1}}n_{1}^{n_{2}}\cdots n_{\ell-1}^{n_{\ell}}\,.$ (S.49)
We here use the relation $n_{j}!\geq(n_{j}/e)^{n_{j}}$ and $(1+y/x)^{x}\leq e^{y}$ for $x>0$ and $y>0$ to get
$\sum_{w_{\rm cl}^{\rm r}}{\|\tilde{h}(\xi_{1},\cdots,\xi_{n})\|\over n!}$ $\leq{(2kgu)^{n}}\sum_{\ell=1}^{n}\sum_{n_{1}+\cdots+n_{\ell}=n\atop n_{i}\geq 1}e^{n_{1}}(|L|/k/n_{1})^{n_{1}}e^{n_{2}}(n_{1}/n_{2})^{n_{2}}\cdots e^{n_{\ell}}(n_{\ell-1}/n_{\ell})^{n_{\ell}}$
$\leq{(2kgue)^{n}}\sum_{\ell=1}^{n}\sum_{n_{1}+\cdots+n_{\ell}=n\atop n_{i}\geq 1}(1+|L|/k/n_{1})^{n_{1}}(1+n_{1}/n_{2})^{n_{2}}\cdots(1+n_{\ell-1}/n_{\ell})^{n_{\ell}}$
$\leq{(2kgue^{2})^{n}}e^{|L|/k}\sum_{\ell=1}^{n}\sum_{n_{1}+\cdots+n_{\ell}=n\atop n_{i}\geq 1}$
$={1\over 2}{(4kgue^{2})^{n}}e^{|L|/k}$
$\leq{1\over 2}(4kgue^{2})^{n}e^{(|X|+|Y|+s^{\prime}k)/k},$ (S.50)
@@
where at the last line we have used the relation $|L|<|X|+|Y|+s^{\prime}k$. At the third line, we use the relation, $\sum_{\ell=1}^{n}\sum_{n_{1}+\cdots+n_{\ell}=n\atop n_{i}\geq 1}=2^{n-1}$. Finally we sum over $s^{\prime}$ to get the following bound.
$\|\rho_{\rm cl}\|_{1}$ $\leq{u^{-1}\over 2}|X|\,|Y|{e^{(|X|+|Y|)/k}\over(1+d_{X,Y})^{\alpha}}\sum_{s=0}^{\infty}(4\beta guke^{2}\,e^{2\beta guk})^{s}\sum_{s^{\prime}=1}^{s}{(1/2ke)^{s^{\prime}}}$
$\leq u^{-1}|X|\,|Y|{e^{(|X|+|Y|)/k}\over(1+d_{X,Y})^{\alpha}}{1\over 1-4\beta guke^{2}\,e^{2\beta guk}}\,.$ (S.51)
@@
Fig. 7(a) illustrates the average mutual information computed over 20,000 random values of $t_{i,j}$ sampled from the interval [0,1][0,1][ 0 , 1 ]. The results reveal that the mutual information increases following the behavior of $N^{2-2\alpha}$ for $\alpha<1$, while saturating when $\alpha>1$.
@@
Figure 7: (a) The mutual information $\mathcal{I}_{\rho_{\beta}}(A:B)$ between subsystem $A$ (the first $N/2$ sites) and subsystem $B$ (the remaining $N/2$ sites) is computed. The system size $N$ varies from 100 to 1000 in increments of 100, while the interaction power $\alpha$ varies from 0.5 to 1.4 in increments of 0.1. The data represents the average mutual information over 20,000 samples, with the hopping parameters $t_{i,j}$ randomly selected from the interval [0,1][0,1][ 0 , 1 ]. For $\alpha<1$, a line proportional to $N^{2-2\alpha}$ is included in the figure for comparison. (b) The average mutual information at $\alpha=1.0$, calculated from 100,000 samples, is plotted as circles with the $x$-axis representing the natural logarithm of the system size, $\ln N$. The red line corresponds to the linear fit, given by $\mathcal{I}_{\rho_{\beta}}(A:B)=0.0575\ln N+0.2082$. (c) The circles (diamonds) represent the exponents $\Delta$, obtained by fitting the average mutual information from panel (a) to a power-law, $\mathcal{I}_{\rho_{\beta}}(A:B)\approx cN^{\Delta}$, over the ranges $N=100$ to $N=500$ (resp. $N=600$ to $N=1000$). The red line represents $\Delta=2-2\alpha$. For $\alpha<1$, the exponent $\Delta$ closely follows the line $2-2\alpha$, whereas for $\alpha>1$, $\Delta$ stays near zero. This behavior becomes more pronounced as the system size increases.
@@
where $a_{i,j}$ is a parameter of order $\mathcal{O}(1)$, $N$ is the number of sites, and ${\bm{S}}_{i}=(S_{i}^{x},S_{i}^{y},S_{i}^{z})$ is a spin-1/2 operator at site $i$. To support our argument that the thermal area holds for $\alpha>(D+1)/2$ and this criterion is optimal, we compute the mutual information $\mathcal{I}_{\rho_{\beta}}(A:B)$ between subsystems $A$ and $B$ at $\beta=2$, with $A$ representing the first $N/2$ sites and $B$ the remaining $N/2$ sites, by using the exact diagonalization method. We randomly assign $a_{i,j}$ from the interval [0,1][0,1][ 0 , 1 ], and then compute the average over 10000 samples. Fig. 8 shows that the mutual information increases with the system size when $\alpha<1$, while it does not exhibit an increase for $\alpha>1$.
Figure 8: Exact diagonalization results: the averaged mutual information $\mathcal{I}_{\rho_{\beta}}(A:B)$ over 10000 samples of the long-range Heisenberg chain (S.53) with random numbers $a_{i,j}$ assigned from the interval [0,1][0,1][ 0 , 1 ]. Here $N$ is the number of sites, $\beta=2$, and $A$ and $B$ are the first $N/2$ sites and the remaining $N/2$ sites, respectively.
@@
We consider $n$ spinless fermionic particles with the basis states |0|0$\rangle$| 0 ⟩ and $|1\rangle=c^{\dagger}|0\rangle$. The determinant of the $n\times n$ matrix $\mathbf{X}=(x_{i,j})$ is defined as